# Symmetry aspects in an ideal Bose gas at thermal equilibrium

### v1

@@ III.1 Broken aggregate phase gauge symmetry
The concept of broken gauge symmetry in solids was originally introduced to understand and describe the occurrence of distinct phases between a critical temperature below which solids undergo a certain phase transition as characterized by an order parameter. In the present theory, such phase transitions in atomic gases are modeled in terms of the composed average wave field $\psi=\psi_{+}\cup\psi_{-}$ that predicts non-zero values below the critical temperature for Bose-Einstein condensation. Different from standard approaches known so far, within the present theory, it is, in particular, possible to explain that it is not the entire gauge field symmetry that is broken, but only the relative symmetry between condensate and non-condensate fields. The two field components differ in terms of an angle of $\pi$, i.e., are directed in the corresponding opposite directions, while the total symmetry of the gas remains preserved, i.e. $\psi=0$.
Please note that this is the case only in three-dimensional setups. For reduced dimensionality, the assumption that $\psi=0$ certainly doesn’t remain valid, since the particle number is not conserved only within manifolds of the total setup, i.e., one has to assume a general function $\psi=\psi(N-N_{0})$. The broken phase gauge symmetry of the (left-shifted) condensate quantum field is illustrated in the lower right section of Fig. 2. As compared to the star-shaped quantum field in Fig. 3, it is visible that the symmetry is broken concerning the zero point axis.
@@ III.3 Direction of time from constructive interference
As we understand from the present model, the direction of time manifests itself as an interference process of forward and backward propagating wave fields $\psi_{+}$ and $\psi_{-}$. Considering only quantum field states with complex number representations within the shown manifold, explains directional time as a propagating quantity that arises from the constructive interference of partial matter waves in the Bose gas. The smallest unit time scales for the forward propagation of time is thus defined by the time scales at which partial matter waves interfere (from zero to a few microseconds), whereas, from destructive interference (dynamical localization in complex space), one may deduce that there exists a certain coherence time of the total fields that defines a time scale for the decay of different realizations of the thermal Boltzmann equilibrium in the present parameter regime.
Figure 7: (Color online) Figure shows $10^{5}$ realizations of the imaginary parts (upper panel - labeled $t_{1}=\Gamma^{-1}_{1}$) and real parts (lower figure - labeled $t_{2}=\Gamma^{-1}_{2}$) of complex phase as defined in Eq. (19) from random Markov sampling for a non-isotropic trap geometry $\omega_{x},\omega_{y}=2\pi\times 42.0$ Hz and $\omega_{z}=120.0$ Hz at temperature $T=5.0$ nK. The observation that $\Gamma_{2}t$ is non-zero at thermal equilibrium indicates the coupling of the equilibrium state to quantum states with $\Gamma_{2}t\neq 0$ (non-classical correlations).
@@
Figure 8: (Color online) Figure shows the coherence phase $t_{2}=\Gamma^{-1}_{2}$ as a function of temperature $T$ for a particle in the trap geometry $\omega_{x},\omega_{y}=2\pi\times 42.0$ Hz and $\omega_{z}=120.0$ Hz. For intermediate values of temperature, the coherence time is around $(300-450)~{}\mu$s. Above the critical temperature, it approaches a value of around $200~{}\mu$s.
@@
Applying only the constraint of particle number conservation as defined by the conservation equation in Eq. (9), numerical sampling leads to a distribution of typical scales for $\Gamma_{1}t$ (oscillation time) and $\Gamma_{2}t$ (coherence time), shown in Fig. 7. From the numerical sampling, it is observed that $\Gamma_{1}t$ is proportional to the real part of the quantum phase and distributed around zero measure in the given parameter range (as defined in Fig. 7 - upper panel). The real part $\Gamma_{1}t$ defines the average oscillation frequency, which means that the quantum particle (with constant particle number at finite temperature) quickly couples to excited single particle quantum states, related to $\Gamma_{1}t\sim 0$ in the form of a standing wave. The distribution of the phase $\Gamma_{2}t$, that is, the effective coherence time for the particle weighted with the decay constant $\Gamma_{2}$, is shown in the lower panel of Fig. 7. The typical range of the time scale $\Gamma^{-1}_{2}$ is between $300~{}\mu s$ and $450~{}\mu s$ (see Figs. 7 and 8) in the present setup. To estimate the time scale in a physical framework, time is scaled in terms of the quantity $\tau=2\pi\times(\omega_{x}\omega_{y}\omega_{z})^{-1/3}$, since any frequency of the equilibrated system below that value is effectively zero in the thermally stable quantum system, because of the finite energy uncertainty. From the present model, it is possible to, in particular, conclude that the so-derived time variable of a bosonic quantum field can provide a fundamental quantization of time only on the scale of quantized time steps $\tau_{0}$ defined by different realizations of the Boltzmann equilibrium. Time intervals smaller than the value of $\tau_{0}$ randomly vary from effectively zero to tens of microseconds at different realizations of the quantum field, which leads to a quasi-continuous numerical time variable without a fundamental unit time for time steps smaller than the time scale $\tau_{0}$.
@@ IV Discussion
In a local representation of the quantum field, in very generality and mathematical and formal terms, the symmetry of the quantum system can formally only be broken from a formal process that either modifies the fundamental equation of state or the corresponding gauge fields, respectively, called gauge fixing ref-16 . In ideal Bose gases, such formal projection can, therefore, only be understood by picking out a certain phase variable by a spooky unphysical action, such as the definition of a mathematical variable by an external user. As shown in the sequel of the present theory, following the ansatz of a non-local order parameter, the occurrence of spontaneously broken gauge fields can be ideally explained in physical terms only assuming the convergence of the (independent) quantum fields to the Boltzmann equilibrium with a vanishing total chemical potential at constant particle number and temperature of the Bose gas. Thus, the symmetry of the considered wave field is spontaneously broken without defining a specific formal gauge of the wave field, but only from the physical constraint of the thermal equilibrium, while the underlying total symmetry of the wave field in complex number representation remains preserved. Finally, the one-directional propagation of time is a direct formal consequence of the symmetry breaking of the Schrödinger equation from the assumption of a Boltzmann equilibrium.
@@ V Conclusion
In summary, one may argue that the presented particle-number conserving quantum field theory explains the concept of broken gauge symmetry as a physical process - the equilibration of the Bose gas to the Boltzmann equilibrium - in weakly interacting Bose-Einstein condensates in the framework of a theory with non-local order parameters. Within the present theory, it was shown that spontaneous symmetry breaking doesn’t rely on defining a specific gauge of the wave field, but only on the implementation of physical constraints in terms of particle number conservation and the definition of a constant temperature that naturally defines a gauge for the wave fields as a solution of the Schrödinger equation in terms of Glauber coherent states. In particular, besides the formal symmetry breaking of the Schrödinger equation’s gauge fields around pre-defined values of zero phase that arises from assuming the Boltzmann equilibrium, i.e. from projecting the field onto a manifold and focussing on the interference of forward and backward propagating components of the wave field, it is possible to identify an intrinsic broken gauge symmetry of the (two) separate components condensate and non-condensate below the critical point for Bose-Einstein condensation. The latter occurs spontaneously below the critical temperature, while the assumption of a thermal Boltzmann equilibrium also leads to one-directional time propagation of the wave field with mainly zero-valued phases above the critical point.
@@ Acknowledgements.
The author acknowledges the financial support from IU Internationale Hochschule for the lecturer position at the university, which has, in particular, enabled the formulation and editing of the present symmetry aspects of ideal Bose gases.

### v2

@@ III.1 Broken aggregate phase gauge symmetry
The concept of broken gauge symmetry in solids was originally introduced to understand and describe the occurrence of distinct phases between a critical temperature below which solids undergo a certain phase transition as characterized by an order parameter. In the present theory, such phase transitions in atomic gases can be modeled by describing the transition of a mixed wave field $\psi=\psi_{0}\cup\psi_{\perp}$ without coherences and a relative reference frame between condensate and non-condensate components of the Bose gas (that leads to non-zero field distribution in the complex plane) to a coherently interfering condensate and non-condensate quantum field that is gauged to zero average (vanishing) values below the critical temperature for Bose-Einstein condensation. In particular, below the critical temperature, the two sub-components condensate and non-condensate are coherently coupled within a certain coherence time (at each realization of the quantum field).
Different from standard approaches known so far, within the present theory, it is, in particular, possible to explain that it is not the entire gauge field symmetry that is broken, but only the relative symmetry between condensate and non-condensate fields that are coherently interacting below the critical temperature. In contrast, in the limit of large temperatures, the two components do not build a relative reference frame for symmetry breaking but are independently mixed, and therefore the total quantum field builds a ball-shaped distribution in complex space. For finite temperatures, there is theoretically no non-zero coherence, however, different field components only interfere destructively, if the coherence time is larger than the time uncertainty. This defines a physical transition for symmetry breaking when the coherence time exceeds this (minimal) uncertainty threshold.
Please note that this is the case only in three-dimensional setups, since in lower dimensions the single-particle wave functions cannot be assumed to obey a non-zero average. Thus, for reduced dimensionality, the assumption that $\psi=0$ doesn’t remain valid, i.e. one has to assume a general function $\psi=\psi(N-N_{0})$. The broken phase gauge symmetry of the (left-shifted) condensate quantum field is illustrated in the lower right section of Fig. 2. As compared to the star-shaped quantum field in Fig. 3, it is visible that the (relative) symmetry is broken concerning the zero point axis assuming thermal equilibrium.
@@ III.3 Direction of time from constructive interference
As we understand from the present model, the direction of time manifests itself as an interference process of forward and backward propagating wave fields $\psi_{+}$ and $\psi_{-}$. Considering only quantum field states with complex number representations within the shown manifold, explains directional time as a propagating quantity that arises from the constructive interference of partial matter waves in the Bose gas. The smallest unit time scales for the forward propagation of time is thus defined by the time scales at which partial matter waves interfere (from zero to a few microseconds), whereas, from destructive interference (dynamical localization in complex space), one may deduce that there exists a certain coherence time of the total field that defines a time scale for the decay of different realizations of the thermal Boltzmann equilibrium in the present parameter regime.
Figure 7: (Color online) Figure shows $10^{5}$ realizations of the imaginary parts (left figure - labeled $t_{1}=\Gamma^{-1}_{1}$) and real parts (right figure - labeled $t_{2}=\Gamma^{-1}_{2}$) of complex phase as defined in Eq. (19) from random Markov sampling for a non-isotropic trap geometry $\omega_{x},\omega_{y}=2\pi\times 42.0$ Hz and $\omega_{z}=120.0$ Hz at temperature $T=5.0$ nK for $N=1000$ atoms. The observation that $\Gamma_{2}t$ is non-zero at thermal equilibrium indicates the coupling of the equilibrium state to quantum states with $\Gamma_{2}t\neq 0$ (non-classical correlations).
@@
Figure 8: (Color online) Figure shows the coherence phase $t_{2}=\Gamma^{-1}_{2}$ as a function of temperature $T$ for a particle in the trap geometry $\omega_{x},\omega_{y}=2\pi\times 42.0$ Hz and $\omega_{z}=120.0$ Hz. For intermediate values of temperature, the coherence time is around $(300-450)~{}\mu$s (larger than the minimal time uncertainty of about $(50-100)~{}\mu$s). Above the critical temperature of about $T_{c}=26.9$ nK, it approaches a value of less than $100~{}\mu$s converging to zero in the limit of very large temperatures.
@@
Applying only the constraint of particle number conservation as defined by the conservation equation in Eq. (9), numerical sampling leads to a distribution of typical scales for $\Gamma_{1}t$ (oscillation time) and $\Gamma_{2}t$ (coherence time), shown in Fig. 7. From the numerical sampling, it is observed that $\Gamma_{1}t$ is proportional to the real part of the quantum phase and distributed around zero measure in the given parameter range (as defined in Fig. 7 - upper figure). The real part $\Gamma_{1}t$ defines the average oscillation frequency, which means that the quantum particle (with constant particle number at finite temperature) quickly couples to excited single particle quantum states, related to $\Gamma_{1}t\sim 0$ in the form of a standing wave. The distribution of the phase $\Gamma_{2}t$, that is, the effective coherence time for the particle weighted with the decay constant $\Gamma_{2}$, is shown in the lower figure of Fig. 7. The typical range of the time scale $\Gamma^{-1}_{2}$ is between $300~{}\mu s$ and $450~{}\mu s$ (see Figs. 7 and 8) in the present setup. To estimate the time scale in a physical framework, time is scaled in terms of the quantity $\tau=2\pi\times(\omega_{x}\omega_{y}\omega_{z})^{-1/3}$, since any frequency of the equilibrated system below that value is effectively zero in the thermally stable quantum system, because of the finite energy uncertainty. From the present model, it is possible to, in particular, conclude that the so-derived time variable of a bosonic quantum field can provide a fundamental quantization of time only on the scale of quantized time steps $\tau_{0}$ defined by different realizations of the Boltzmann equilibrium. Time intervals smaller than the value of $\tau_{0}$ randomly vary from effectively zero to tens of microseconds at different realizations of the quantum field, which leads to a quasi-continuous numerical time variable without a fundamental unit time (i.e. varying time step sizes) for time steps smaller than the time scale $\tau_{0}$.
@@ IV Discussion
In a local representation of the quantum field, in very generality and mathematical and formal terms, the local symmetry of the quantum system can formally only be broken from a formal process that either modifies the fundamental equation of state or the corresponding gauge fields, respectively, called gauge fixing ref-16 . In ideal Bose gases, such formal projection can, therefore, not only be understood by picking out a certain phase variable by a spooky unphysical action, such as the definition of a mathematical variable by an external observer. As shown in the sequel of the present theory, following the ansatz of a non-local order parameter, the occurrence of spontaneously broken gauge fields can be ideally explained in physical terms by assuming the convergence of the (independent) quantum fields to the Boltzmann equilibrium with a vanishing total chemical potential at constant particle number and temperature of the Bose gas. Thus, the symmetry of the considered wave field is spontaneously broken not from defining a specific formal (zero) gauge of the wave field, but from the physical inset of quantum coherence at the thermal equilibrium, while the underlying total symmetry of the wave field in complex number representation remains preserved. Finally, the one-directional propagation of time is a direct formal consequence of the symmetry breaking of the Schrödinger equation from the assumption of a Boltzmann equilibrium.
@@ V Conclusion
In summary, the presented particle-number conserving quantum field theory explains the concept of broken gauge symmetry as a physical rather than a purely formal and mathematical process from coherent interference of cold atomic matter waves and the equilibration of the Bose gas to the Boltzmann equilibrium. Within the present theory, it was shown that spontaneous symmetry breaking doesn’t rely only on a formal definition of a specific gauge of the wave field, but that symmetry breaking with a certain zero gauge of the wave field and one-directional time propagation naturally follows from the implementation of physical constraints in terms of particle number conservation and the assumption of the Boltzmann equilibrium. This is because, below the critical temperature, the manifolds of condensate and non-condensate quantum fields interfere destructively as a consequence of the increasing coherence time between different single-particle states and define a relative reference frame for the relative symmetry breaking between condensate and non-condensate aggregate phases in the Bose gas.
@@ Acknowledgements.
The author acknowledges the financial support from IU Internationale Hochschule for the lecturer position at the university, which has, in particular, enabled the formulation and editing of the present theory on symmetry aspects of ideal Bose gases.